OpenStudy (anonymous):

Basic Complex Variables question. Can someone please explain to me how this : (x1, y1)(x2, y2) = (x1x2 − y1y2, y1x2 + x1y2). Is come to?

OpenStudy (helder_edwin):

it is only a definition. it is the definition of multiplication of complex numbers.

OpenStudy (anonymous):

I understand that, but I dont understand the process in which you go through. could an example be given?

OpenStudy (anonymous):

are those x^2 and y^2 or x1y2?

OpenStudy (anonymous):

x sub 1, x sub 2, etc...

OpenStudy (helder_edwin):

right. \[ \large (2,3)\cdot(1,5)=(2\cdot1-3\cdot5,2\cdot5+3\cdot1) \] \[ \large =(2-15,10+3)=(-13,13) \]

OpenStudy (anonymous):

Thank you. magepker728, you can't distribute these. They are coordinates not binomials.